# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ 4.3 Discussion: inferred parameters
4.4 Discussion: the predicted VHE emission of GW 170817
We now discuss the predicted VHE emission of the afterglow of GW 170817. The predicted light curve at 1 TeV is plotted in Fig. 8 and the VHE component of the spectrum at the peak of the light curve is shown in Fig. 9, where is also indicated the upper limit obtained by H.E.S.S. (Abdalla et al., 2020). It appears that (i) the synchrotron emission extends at most up to the GeV range [^7] The emission in the $0.1-1\,\mathrm{GeV}$ range is then dominated by the synchrotron radiation. For comparison with Fig. 9, the sensitivity of Fermi/LAT above 100 MeV is of the order of $10^{-11}\,\mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}$ for a $\sim 1\,\mathrm{day}$ observation (Abdollahi et al., 2020, 2022). The afterglow of GW 170817 was therefore undetectable by Fermi/LAT at any time. (due to the synchrotron burnoff limit, see Appendix A) and the VHE emission is then entirely due to the SSC process; (ii) even taking into account the dispersion, the predicted flux at 1 TeV is at least two orders of magnitude below the upper limit of $3.2\times 10^{-13}\leavevmode\nobreak\ \mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}$ obtained around the peak by H.E.S.S. (Abdalla et al., 2020), which is therefore not constraining. It is also clear when comparing Figs. 7 and 8 that the dispersion of the predicted VHE flux is much larger than for the synchrotron component which is well constrained by the observations from radio to X-rays. This is also shown in the distribution of the predicted peak flux and peak time in radio and at 1 TeV in Fig. 10. Note that the VHE light curve peaks a few days before the light curves from radio to X-rays, due to a different evolution of the synchrotron and SSC components. This result is due to the fact that the dependency of the SSC component on the model parameters is complex and very different from the dependencies of the synchrotron component identified by Sari et al. (1998). This also highlights the interest of detecting the VHE emission for the inference of model parameters, as some of the degeneracies visible in Fig. 6 would be broken.
In the afterglow of GW 170817, the synchrotron spectrum is in slow cooling, with a spectral peak determined by the critical Lorentz factor $\gamma_{\mathrm{c}}$. If IC scatterings occurred in Thomson regime, the peak of the SSC component would be due to the diffusion of photons at $\nu_{\mathrm{c}}$ by the electrons at $\gamma_{\mathrm{c}}$ (Sari & Esin, 2001), and this peak should be intense as $\epsilon_{\mathrm{e}}\gg\epsilon_{\mathrm{B}}$ (see Equation (77)). However, the predicted VHE emission is actually very weak because of a strong KN attenuation. To illustrate this effect, we select in Table 8 two reference sets of parameters, taken from the posterior sample. They are respectively in the 68% and 97.5% confidence intervals for the ”SSC (with KN)” fit and fit perfectly well the observed light curves. The first set of parameters (”Moderate”) corresponds to a predicted VHE flux which is close to the median, whereas the second set of parameters (”Optimistic”) corresponds to a predicted VHE flux on the highest end of the confidence interval: see the two predicted VHE light curves and peak spectra in Fig. 11. In both cases, we selected sets of parameters having a low density, and therefore a low energy (see discussion in § 4.3).
@@ 4.4 Discussion: the predicted VHE emission of GW 170817
Parameter ”Moderate” ”Optimistic”
@@
888 Both sets of parameters are in the 97.5% confidence interval of the distributions inferred by the ”SSC (with KN)” fit of the afterglow of GW 170817. They differ by the predicted VHE flux, which is either close to the median (”Moderate” case) or at the higher end of the confidence interval (”Optimistic” case).
@@
Figure 11: VHE light curve at $h\nu_{\mathrm{obs}}=1\leavevmode\nobreak\ \mathrm{TeV}$ (top) and spectrum around the peak at $t_{\mathrm{obs}}=110\leavevmode\nobreak\ \mathrm{days}$ (bottom) for the ”Moderate” (in blue) and ”Optimistic” (in red) references cases (see Table 8). Both predict similar fluxes in the synchrotron range due to the observational constraints, but differ in the SSC regime. The upper limit from H.E.S.S. is indicated for comparison (Abdalla et al., 2020).
@@
Fig. 12 shows for these two reference cases the evolution of $Y(\gamma_{\mathrm{c}})$ and $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}$ for the core jet, computing the SSC emission either by assuming that all scatterings are in Thomson regime like in the ”SSC (Thomson)” fit, or by including the KN attenuation like in the ”SSC (with KN)” fit. The figure shows that, in the full calculation, $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}\gg 1$ during the whole evolution, so that the scatterings of photons at $\nu_{\mathrm{c}}$ by electrons at $\gamma_{\mathrm{c}}$ are strongly reduced by the KN regime. The Compton parameter $Y(\gamma_{\mathrm{c}})$ then remains very low. This confirms the origin of the weak VHE emission in the afterglow of GW 170817. In this regime of weak SSC emission, the synchrotron component is unaffected by the IC cooling, $\gamma_{\mathrm{c}}\simeq\gamma_{\mathrm{c}}^{\mathrm{syn}}$, and this explains why the results of the ”no SSC” and ”SSC (with KN)” fits are so close (Table 5). On the other hand, Fig. 12 also shows that $Y^{\mathrm{no\,KN}}$ overestimates $Y(\gamma_{\mathrm{c}})$ in this case, leading to $Y(\gamma_{\mathrm{c}})>1$. The critical Lorentz factor is therefore given by $\gamma_{\mathrm{c}}\simeq\gamma_{\mathrm{c}}^{\mathrm{syn}}/Y(\gamma_{\mathrm{c}})$, which leads to an overestimation of $\gamma_{\mathrm{c}}^{\mathrm{syn}}$ in the MCMC to maintain the cooling break above the X-ray measurements (increased $E_{0\mathrm{,iso}}$ and lower $\epsilon_{\mathrm{B}}$). This explains the observed difference between the values of the model parameters inferred in the ”SSC (Thomson)” and ”SSC (with KN)” fits in Table 5. The value of $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}\gg 1$ in this case (green curves in the top panel of Fig. 12) proves that the Thomson regime is not justified.
@@
As the upper limit of H.E.S.S. was obtained close to the predicted peak with an already long exposure time of 53.9 hours (Abdalla et al., 2020), our prediction shows that it was unfortunately impossible to detect the VHE afterglow of GW 170817 with instruments currently available. We discuss in the next section the conditions that would make a post-merger afterglow detectable at VHE in the future, especially in the context of the increased sensitivity expected with the CTA.
@@
Figure 12: Evolution of $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}$ (top) and of the Compton parameter at $\gamma_{\mathrm{c}}$ , $Y(\gamma_{\mathrm{c}})$ (bottom) in the core jet. These quantities are plotted as a function of the observer time for the ”Moderate” (solid line) and ”Optimistic” (dashed line) references cases (see Table 8) in the ”SSC (Thomson)” case (green) and in the ”SSC (with KN)” case (orange).
@@
5 Detectability of post-merger afterglows at VHE and prospects for the CTA
@@
Even if the VHE afterglow of GW 170817 was not detectable, deeper observations of similar events can be expected in the future, in particular with the Cerenkov Telescope Array (CTA, Cherenkov Telescope Array Consortium et al. 2019). The off-axis afterglow phase is of particular interest because after a few days, full-night observations can be conducted without a significant intrinsic source flux variability over the observation time. We investigate in this section the conditions that would make such an event detectable by current instruments and the upcoming CTA. For this study, we assume a CTA detection sensitivity at $1\leavevmode\nobreak\ \mathrm{TeV}$ about 6 times better than that of H.E.S.S., as expected from Cherenkov Telescope Array Consortium et al. (2019). For an exposure time of $\sim 50\,\mathrm{hours}$ as for the H.E.S.S. deep observations of GW 170817 (Abdalla et al., 2020), this leads to a detection sensitivity at $1\leavevmode\nobreak\ \mathrm{TeV}$ of $5\times 10^{-14}\leavevmode\nobreak\ \mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}$.
@@
A geometrical effect such as a lower viewing angle naturally increases the peak of the observed flux, and has a similar effect on the synchrotron and SSC components. In Fig. 13, we show for the ”Moderate” reference case the light curves (top) and spectra at the time of the $1\leavevmode\nobreak\ \mathrm{TeV}$ emission peak (bottom), for gradually decreasing viewing angles. The $1\leavevmode\nobreak\ \mathrm{TeV}$ light curves peak at earlier times when the viewing angle decreases, and the peak flux reaches higher values. As expected, the relative flux increase at the time of peak is similar for the synchrotron and the SSC components (Fig. 13, bottom) and all light curves follow a similar long-term evolution, when the entire jet emission becomes visible to the off-axis observer. For the ”Moderate” (resp. ”Optimistic”) reference case (Table 8), we find that a ratio $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\lesssim 3$ (resp. $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\lesssim 6$) would have made the afterglow of GW 170817 detectable by H.E.S.S., provided the conditions had allowed for a very early deep observation. With the CTA, the VHE emission could have been detected up to $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\simeq 4$ (resp. $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\simeq 8$). In such a scenario, the emission peaks at $\sim 5\leavevmode\nobreak\ \mathrm{days}$ (resp. $13\leavevmode\nobreak\ \mathrm{days}$). This opens up interesting perspectives, even if the expected detection rate of nearby events at small viewing angles is low (see e.g. Duque et al., 2019; Mochkovitch et al., 2021).
@@
Under some conditions, the predicted VHE emission could also be intrinsically brighter. As discussed in § 4.4, the VHE afterglow of GW 170817 is weak because of a strong KN attenuation ($\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}\gg 1$). The regime for the diffusions of photons at $\nu_{\mathrm{c}}$ by electrons at $\gamma_{\mathrm{c}}$ is affected by several parameters (Nakar, 2007). When the diffusions occur in the Thomson regime, the predicted scaling in the core jet is
@@
Therefore, the Thomson regime is favored by a high magnetic field, a high density or a high kinetic energy. As $\epsilon_{\mathrm{B}}$ is determined by plasma instabilities at the ultra-relativistic shock (Sironi et al., 2015), it should not vary much from an afterglow to another. On the other hand, the external density and the kinetic energy may differ between mergers. We focus here on the effect of the external density, which is slightly stronger. Fig. 14 shows the evolution of the afterglow emission for the ”Moderate” reference case when $n_{\mathrm{ext}}$ increases up to $\sim 10\,\mathrm{cm^{-3}}$, all other parameters being kept constant. A higher external density leads to an earlier peak time at higher luminosities but affects differently the synchrotron and the SSC flux: as expected the SSC process becomes more efficient and the relative SSC-to-synchrotron ratio increases. The transition from the weak SSC emission due to a strong KN attenuation described in § 4.4 to a more efficient case in Thomson regime occurs in this reference case above $n_{\mathrm{ext}}\sim 1\leavevmode\nobreak\ \mathrm{cm}^{-3}$. This corresponds to the radiative regime S15 in Appendix B, with spectral breaks at $\nu_{\mathrm{m}}<\nu_{\mathrm{c}}<\widehat{\nu}_{\mathrm{c}}<\nu_{0}$. This effect can clearly be observed on the bottom panel of Fig. 14, showing the observed spectrum at the peak emission. At the largest external densities, the increase of the VHE flux stalls for two reasons. Firstly, the strong SSC emission starts to affect significantly the synchrotron spectrum, which features two peaks (visible on the bottom panel of Fig. 14). The cooling frequency $\nu_{\mathrm{c}}$ at the peak also decreases, shifting the SSC peak to lower frequencies so that the flux at 1 TeV does not increase as significantly as the SSC peak flux. A second limitation, dominant at the highest densities, comes from the pair production.
@@
Even taking into account these limitations, a higher external density clearly favors a brighter VHE emission, even if both the ”Moderate” and ”Optimistic” reference cases remain undetectable by H.E.S.S. or the CTA at the reference viewing angle. We conclude that post-merger VHE afterglows should become detectable in the future only under several favourable conditions, e.g. a higher external density and a slightly lower viewing angle. For instance, the ”Moderate” (resp. ”Optimistic”) reference case with an external density $n_{\mathrm{ext}}=1\leavevmode\nobreak\ \mathrm{cm^{-3}}$ and a viewing angle $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}=6$ should be detectable by the CTA [^9] On the other hand, even with $n_{\mathrm{ext}}=1\,\mathrm{cm^{-3}}$ and $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}=6$, the emission in the Fermi/LAT range at the peak is still dominated by the synchrotron radiation and remains undetectable at $40\,\mathrm{Mpc}$ for a $\sim 1\,\mathrm{day}$ observation. up to a distance of $\sim 90\leavevmode\nobreak\ \mathrm{Mpc}$ (resp. $\sim 400\leavevmode\nobreak\ \mathrm{Mpc}$). Note however that two additional effects should be included in the ”Optimistic” case and may reduce this maximum distance: (i) at 400 Mpc, the attenuation by the extragalactic background light is not negligible anymore at 1 TeV (see e.g. Domínguez et al. 2011); (ii) in this case, the VHE afterglow peaks very early ($\sim 1$ day). This would require an early follow-up which will necessarily be less deep due to a limited exposure time. It remains that detections with the CTA above 100 Mpc are possible in the ”Optimistic” case. As some parameters that are naturally degenerate in an afterglow fit including only synchrotron radiation, such as $n_{\mathrm{ext}}$ and $\epsilon_{\mathrm{B}}$, have different effects on the SSC emission, such VHE detections would therefore better constrain the parameter space and allow for a better understanding of the physical conditions in the post-merger relativistic ejecta.
Figure 13: Light curves at $1\leavevmode\nobreak\ \mathrm{TeV}$ (top panel) and spectra at light curve peak (bottom panel) of the ”Moderate” reference case (see Table 8) for varying viewing angles. The case fitting the afterglow of GW 170817 corresponds to $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}=12.17$. The peak times at which spectra are calculated are (in the ordering of the legend) $123$, $73$, $39$, $17$, $4.7$, $1.4$ and $0.06$ days. The dotted line in the top panel corresponds to the assumed CTA sensitivity in Section 5.
@@
Figure 14: Light curves at $1\leavevmode\nobreak\ \mathrm{TeV}$ (top panel) and spectra at light curve peak (bottom panel) of the ”Moderate” reference case (see Table 8) for varying external medium densities $n_{\mathrm{ext}}$. The case fitting the afterglow of GW 170817 corresponds to $n_{\mathrm{ext}}=4.25\times 10^{-3}\leavevmode\nobreak\ \mathrm{cm}^{-3}$. Each line is labeled with the multiplication factor applied to the reference $n_{\mathrm{ext}}$. The peak times at which spectra are calculated are $123$, $60$, $26$, $12$ and $0.55$ days. The dotted line in the top panel corresponds to the assumed CTA sensitivity in Section 5.

### v2

@@ 4.3 Discussion: Inferred parameters
4.4 Discussion: Predicted very high energy emission of GW 170817
We now discuss the predicted VHE emission of the afterglow of GW 170817. The predicted light curve at 1 TeV is plotted in Fig. 8, and the VHE component of the spectrum at the peak of the light curve is shown in Fig. 9, where we also indicate the upper limit obtained by H.E.S.S. (Abdalla et al. 2020). It appears that (i) the synchrotron emission extends at most up to the GeV range [^7] The emission in the $0.1-1\,\mathrm{GeV}$ range is then dominated by the synchrotron radiation. For comparison with Fig. 9, the sensitivity of Fermi/LAT above 100 MeV is of the order of $10^{-11}\,\mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}$ for a $\sim 1\,\mathrm{day}$ observation (Abdollahi et al. 2020, 2022). The afterglow of GW 170817 was therefore undetectable by Fermi/LAT at any time. (due to the synchrotron burnoff limit; see Appendix A) and the VHE emission is then entirely due to the SSC process; (ii) even when the dispersion is taken into account, the predicted flux at 1 TeV is at least two orders of magnitude below the upper limit of $3.2\times 10^{-13}\leavevmode\nobreak\ \mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}$ obtained around the peak by H.E.S.S. (Abdalla et al. 2020), which is therefore not constraining. It is also clear when comparing Figs. 7 and 8 that the dispersion of the predicted VHE flux is much larger than for the synchrotron component, which is well constrained by the observations from radio to X-rays. This is also shown in the distribution of the predicted peak flux and peak time in radio and at 1 TeV in Fig. 10. The VHE light curve peaks a few days before the light curves from radio to X-rays because the synchrotron and SSC components evolve differently. This result is due to the fact that the dependence of the SSC component on the model parameters is complex and very different from the dependences of the synchrotron component identified by Sari et al. (1998). This also highlights the interest of detecting the VHE emission for the inference of model parameters, as some of the degeneracies visible in Fig. 6 would be broken.
@@ 4.4 Discussion: Predicted very high energy emission of GW 170817
Parameter Moderate Optimistic
@@
888 The two parameter sets are in the 97.5% confidence interval of the distributions inferred by the ”SSC (with KN)” fit of the afterglow of GW 170817. They differ by the predicted VHE flux, which is either close to the median (moderate case) or at the higher end of the confidence interval (optimistic case).
In the afterglow of GW 170817, the synchrotron spectrum is in the slow-cooling regime, with a spectral peak determined by the critical Lorentz factor $\gamma_{\mathrm{c}}$. If IC scattering occurred in Thomson regime, the peak of the SSC component would be due to the scattering of photons at $\nu_{\mathrm{c}}$ by the electrons at $\gamma_{\mathrm{c}}$ (Sari & Esin 2001), and this peak should be intense as $\epsilon_{\mathrm{e}}\gg\epsilon_{\mathrm{B}}$ (see Eq. (77)). However, the predicted VHE emission is very weak because the KN attenuation is strong. To illustrate this effect, we select in Table 8 two reference sets of parameters that are taken from the posterior sample. They are in the 68% and 97.5% confidence intervals for the ”SSC (with KN)” fit and fit the observed light curves perfectly well. The first set of parameters (moderate) corresponds to a predicted VHE flux that is close to the median, whereas the second set of parameters (optimistic) corresponds to a predicted VHE flux at the highest end of the confidence interval (see the two predicted VHE light curves and peak spectra in Fig. 11). In both cases, we selected parameter sets with a low density and therefore, with a low energy (see the discussion in Sect. 4.3).
@@
Figure 11: VHE light curve at $h\nu_{\mathrm{obs}}=1\leavevmode\nobreak\ \mathrm{TeV}$ (top) and spectrum around the peak at $t_{\mathrm{obs}}=110\leavevmode\nobreak\ \mathrm{days}$ (bottom) for the moderate (in blue) and optimistic (in red) reference cases (see Table 8). Both predict similar fluxes in the synchrotron range due to the observational constraints, but they differ in the SSC regime. The upper limit from H.E.S.S. is indicated for comparison (Abdalla et al. 2020).
@@
Fig. 12 shows for these two reference cases the evolution of $Y(\gamma_{\mathrm{c}})$ and $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}$ for the core jet. The SSC emission is either computed by assuming that all scatterings are in Thomson regime, as in the ”SSC (Thomson)” fit, or by including the KN attenuation, as in the ”SSC (with KN)” fit. The figure shows that in the full calculation, $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}\gg 1$ during the whole evolution, so that the scattering of photons at $\nu_{\mathrm{c}}$ by electrons at $\gamma_{\mathrm{c}}$ is strongly reduced in the KN regime. The Compton parameter $Y(\gamma_{\mathrm{c}})$ then remains very low. This confirms the origin of the weak VHE emission in the afterglow of GW 170817. In this regime of weak SSC emission, the synchrotron component is not affected by the IC cooling, $\gamma_{\mathrm{c}}\simeq\gamma_{\mathrm{c}}^{\mathrm{syn}}$, and this explains why the results of the ”no SSC” and ”SSC (with KN)” fits are so close (Table 5). On the other hand, Fig. 12 also shows that $Y^{\mathrm{no\,KN}}$ overestimates $Y(\gamma_{\mathrm{c}})$ in this case, leading to $Y(\gamma_{\mathrm{c}})>1$. The critical Lorentz factor is therefore given by $\gamma_{\mathrm{c}}\simeq\gamma_{\mathrm{c}}^{\mathrm{syn}}/Y(\gamma_{\mathrm{c}})$, which leads to an overestimation of $\gamma_{\mathrm{c}}^{\mathrm{syn}}$ in the MCMC to maintain the cooling break above the X-ray measurements (increased $E_{0\mathrm{,iso}}$ and lower $\epsilon_{\mathrm{B}}$). This explains the observed difference between the values of the model parameters inferred in the ”SSC (Thomson)” and ”SSC (with KN)” fits in Table 5. The value of $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}\gg 1$ in this case (green curves in the top panel of Fig. 12) proves that the Thomson regime is not justified.
@@
As the upper limit of H.E.S.S. was obtained close to the predicted peak with an already long exposure time of 53.9 hours (Abdalla et al. 2020), our prediction shows that it was unfortunately impossible to detect the VHE afterglow of GW 170817 with currently available instruments. We discuss the conditions that would make a post-merger afterglow detectable at VHE in the future in the next section, especially in the context of the increased sensitivity expected with the CTA.
@@
Figure 12: Evolution of $\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}$ (top) and of the Compton parameter at $\gamma_{\mathrm{c}}$ , $Y(\gamma_{\mathrm{c}})$ (bottom) in the core jet. These quantities are plotted as a function of the observer time for the moderate (solid line) and optimistic (dashed line) reference cases (see Table 8) in the ”SSC (Thomson)” case (green) and in the ”SSC (with KN)” case (orange).
@@
5 Detectability of post-merger afterglows at very high energy, and prospects for the CTA
@@
Even if the VHE afterglow of GW 170817 was not detectable, deeper observations of similar events can be expected in the future, in particular, with the CTA. The off-axis afterglow phase is of particular interest because after a few days, full-night observations can be conducted without a significant intrinsic source flux variability over the observation time. We investigate in this section the conditions that would make such an event detectable by current instruments and the upcoming CTA. For this study, we assumed a CTA detection sensitivity at $1\leavevmode\nobreak\ \mathrm{TeV}$ about six times better than that of H.E.S.S., as expected from Cherenkov Telescope Array Consortium et al. (2019). For an exposure time of $\sim 50\,\mathrm{hours}$ as for the H.E.S.S. deep observations of GW 170817 (Abdalla et al. 2020), this leads to a detection sensitivity at $1\leavevmode\nobreak\ \mathrm{TeV}$ of $5\times 10^{-14}\leavevmode\nobreak\ \mathrm{erg}\cdot\mathrm{s}^{-1}\cdot\mathrm{cm}^{-2}$.
Figure 13: Light curves at $1\leavevmode\nobreak\ \mathrm{TeV}$ (top panel) and spectra at light-curve peak (bottom panel) of the moderate reference case (see Table 8) for varying viewing angles. The case fitting the afterglow of GW 170817 corresponds to $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}=12.17$. The peak times at which spectra are calculated are (in the ordering of the legend) $123$, $73$, $39$, $17$, $4.7$, $1.4$, and $0.06$ days. The dotted line in the top panel corresponds to the assumed CTA sensitivity in Sect. 5.
@@
A geometrical effect such as a lower viewing angle naturally increases the peak of the observed flux, and it has a similar effect on the synchrotron and SSC components. In Fig. 13, we show the light curves (top) and spectra at the time of the $1\leavevmode\nobreak\ \mathrm{TeV}$ emission peak (bottom) for the moderate reference case, and for gradually decreasing viewing angles. The $1\leavevmode\nobreak\ \mathrm{TeV}$ light curves peak at earlier times when the viewing angle decreases, and the peak flux reaches higher values. As expected, the relative flux increase at the time of peak is similar for the synchrotron and the SSC components (Fig. 13, bottom), and all light curves follow a similar long-term evolution when the entire jet emission becomes visible to the off-axis observer. For the moderate (optimistic) reference case (Table 8), we find that a ratio $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\lesssim 3$ ($\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\lesssim 6$) would have made the afterglow of GW 170817 detectable by H.E.S.S., provided the conditions had allowed for a very early deep observation. With the CTA, the VHE emission could have been detected up to $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\simeq 4$ ($\theta_{\mathrm{v}}/\theta_{\mathrm{c}}\simeq 8$). In this scenario, the emission peaks at $\sim 5\leavevmode\nobreak\ \mathrm{days}$ ($13\leavevmode\nobreak\ \mathrm{days}$). This opens up interesting perspectives even when the expected detection rate of nearby events at small viewing angles is low (see e.g. Duque et al. 2019; Mochkovitch et al. 2021).
@@
Under some conditions, the predicted VHE emission could also be intrinsically brighter. As discussed in Sect. 4.4, the VHE afterglow of GW 170817 is weak because of the strong KN attenuation ($\gamma_{\mathrm{c}}/\widehat{\gamma}_{\mathrm{c}}\gg 1$). The regime for the scattering of photons at $\nu_{\mathrm{c}}$ by electrons at $\gamma_{\mathrm{c}}$ is affected by several parameters (Nakar 2007). When the scattering occurs in the Thomson regime, the predicted scaling in the core jet is
@@
Therefore, the Thomson regime is favoured by a high magnetic field, a high density, or a high kinetic energy. As $\epsilon_{\mathrm{B}}$ is determined by plasma instabilities at the ultra-relativistic shock (Sironi et al. 2015), it should not vary much from an afterglow to the next. On the other hand, the external density and the kinetic energy may differ between mergers. We focus here on the effect of the external density, which is slightly stronger. Fig. 14 shows the evolution of the afterglow emission for the moderate reference case when $n_{\mathrm{ext}}$ increases up to $\sim 10\,\mathrm{cm^{-3}}$, all other parameters being kept constant. A higher external density leads to an earlier peak time at higher luminosities, but affects the synchrotron and the SSC flux differently: As expected, the SSC process becomes more efficient and the relative SSC-to-synchrotron ratio increases. The transition from the weak SSC emission due to a strong KN attenuation described in Sect. 4.4 to a more efficient case in Thomson regime occurs in this reference case above $n_{\mathrm{ext}}\sim 1\leavevmode\nobreak\ \mathrm{cm}^{-3}$. This corresponds to the radiative regime S15 in Appendix B, with spectral breaks at $\nu_{\mathrm{m}}<\nu_{\mathrm{c}}<\widehat{\nu}_{\mathrm{c}}<\nu_{0}$. This effect can clearly be observed in the bottom panel of Fig. 14, showing the observed spectrum at the peak emission. At the highest external densities, the increase in the VHE flux stalls for two reasons. Firstly, the strong SSC emission starts to affect the synchrotron spectrum significantly, which features two peaks (visible on the bottom panel of Fig. 14). The cooling frequency $\nu_{\mathrm{c}}$ at the peak also decreases, shifting the SSC peak to lower frequencies so that the flux at 1 TeV does not increase as significantly as the SSC peak flux. A second limitation, dominant at the highest densities, comes from the pair production.
@@
Even when these limitations are taken into account, a higher external density clearly favours a brighter VHE emission, even when both the moderate and optimistic reference cases remain undetectable by H.E.S.S. or the CTA at the reference viewing angle. We conclude that post-merger VHE afterglows should become detectable in the future only under several favourable conditions, such as a higher external density and a slightly lower viewing angle. For instance, the moderate (optimistic) reference case with an external density $n_{\mathrm{ext}}=1\leavevmode\nobreak\ \mathrm{cm^{-3}}$ and a viewing angle $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}=6$ should be detectable by the CTA [^9] On the other hand, even with $n_{\mathrm{ext}}=1\,\mathrm{cm^{-3}}$ and $\theta_{\mathrm{v}}/\theta_{\mathrm{c}}=6$, the emission in the Fermi/LAT range at the peak is still dominated by the synchrotron radiation and remains undetectable at $40\,\mathrm{Mpc}$ for a $\sim 1\,\mathrm{day}$ observation. up to a distance of $\sim 90\leavevmode\nobreak\ \mathrm{Mpc}$ ($\sim 400\leavevmode\nobreak\ \mathrm{Mpc}$). However, two additional effects should be included in the optimistic case and may reduce this maximum distance: (i) At 400 Mpc, the attenuation by the extragalactic background light is no longer negligible at 1 TeV (see e.g. Domínguez et al. 2011), and (ii) in this case, the VHE afterglow peaks very early ($\sim 1$ day). This would require an early follow-up, which will necessarily be less deep due to a limited exposure time. Detections with the CTA above 100 Mpc remain possible in the optimistic case. As some parameters that are naturally degenerate in an afterglow fit including only synchrotron radiation, such as $n_{\mathrm{ext}}$ and $\epsilon_{\mathrm{B}}$, have different effects on the SSC emission, these VHE detections would therefore better constrain the parameter space and allow for a better understanding of the physical conditions in the post-merger relativistic ejecta.
@@
Figure 14: Light curves at $1\leavevmode\nobreak\ \mathrm{TeV}$ (top panel) and spectra at light-curve peak (bottom panel) of the moderate reference case (see Table 8) for varying external medium densities $n_{\mathrm{ext}}$. The case fitting the afterglow of GW 170817 corresponds to $n_{\mathrm{ext}}=4.25\times 10^{-3}\leavevmode\nobreak\ \mathrm{cm}^{-3}$. Each line is labelled with the multiplication factor applied to the reference $n_{\mathrm{ext}}$. The peak times at which spectra are calculated are $123$, $60$, $26$, $12$, and $0.55$ days. The dotted line in the top panel corresponds to the assumed CTA sensitivity in Sect. 5.